OpenStudy (anonymous):

The formula I have is to find weight. If I have ther weight(lbs) how do I write this to find Thickness if I have Width, length, and weight Width x Length x Mil ÷ 15 = Lbs./

OpenStudy (anonymous):

I think you're being asked to use 'density'. The density of an object is its mass per unit volume;\[\rho = \frac{m}{V}\]The volume here is (according to you), \[V=(length)\times(width)\times(thickness)\]and you have the mass (you call it weight, but it's mass). You say you want the 'thickness', so that would be"\[\rho = \frac{m}{lwt }\rightarrow t=\frac{m}{\rho lw}\]where l is length w is width t is thickness m is mass p is density.